OpenStudy (anonymous):

Jimmy's toy car is powered by four 1.5 volt batteries. The motor has a resistance of 40 ohms. Yesterday, Jimmy played with the car for 2 hours. How much energy did the batteries provide during that time? A) 1.8J B) 27J C) 108J D) 6480J

OpenStudy (kropot72):

Assume the 4 cells are connected in series. \[P(watts)=\frac{V ^{2}}{R}=\frac{6^{2}}{40}\] Joule = watt-second = \[\frac{6^{2}}{40}\times 2\times 3600=? J\]